# Explainable Benchmarking for Iterative Optimization Heuristics

### 31 Jan 2024 v1

@@ 4.5. Performance comparison between ModCMA and ModDE
While we have to stress that it is not the main aim of this work to compare ModCMA with ModDE, we can do a (limited) comparison in performance between the two modular frameworks using the data collected. In Table 9 we listed the performance of the single-best, avg-best and “default” performance of the frameworks in $5$ dimensions and in Table 10 for $30$ dimensions. The “default” performance denotes the average performance over all algorithm configurations. Results in boldface indicates a significant improvement (with p-value threshold of $0.05$ between the two frameworks in the same category (so single-best vs single-best, avg-best vs avg-best etc.). It can be observed that in general ModCMA seems to be better performing on most of the BBOB functions except for $f_{2},f_{3}$ and $f_{4}$ (all from the separable functions group). However, take into account that not all ModDE and not all ModCMA parameter options are tested (as this is computationally infeasible). If we look at the stability or tune-ability of the different frameworks we can conclude that in general ModDE gains more performance on average by optimizing the hyper-parameters. For ModCMA in 30 dimensions, the gain (in terms of AOCC) to optimize for a specific function is on average not significant over taking the average best configuration (See Table 4.
@@
f1 Sphere 0.97 (0.00) 0.98 (0.00) 0.93 (0.00) 0.97 (0.00) 0.66 (0.26) 0.74 (0.30)
f2 Ellipsoid 0.94 (0.01) 0.91 (0.01) 0.88 (0.01) 0.90 (0.00) 0.48 (0.33) 0.47 (0.37)
f3 Rastrigin 0.89 (0.01) 0.47 (0.41) 0.66 (0.21) 0.17 (0.01) 0.29 (0.25) 0.12 (0.09)
f4 BuecheRastrigin 0.82 (0.03) 0.29 (0.34) 0.46 (0.27) 0.13 (0.01) 0.22 (0.20) 0.09 (0.04)
f5 LinearSlope 1.00 (0.00) 1.00 (0.00) 0.99 (0.01) 0.99 (0.00) 0.94 (0.17) 0.96 (0.14)
f6 AttractiveSector 0.88 (0.05) 0.95 (0.01) 0.75 (0.26) 0.91 (0.00) 0.30 (0.22) 0.54 (0.36)
f7 StepEllipsoid 0.91 (0.01) 0.96 (0.00) 0.89 (0.02) 0.95 (0.02) 0.37 (0.25) 0.45 (0.33)
f8 Rosenbrock 0.81 (0.06) 0.92 (0.01) 0.71 (0.23) 0.86 (0.01) 0.25 (0.21) 0.49 (0.36)
f9 RosenbrockRotated 0.80 (0.03) 0.92 (0.01) 0.70 (0.24) 0.87 (0.01) 0.23 (0.21) 0.52 (0.34)
f10 EllipsoidRotated 0.78 (0.03) 0.91 (0.01) 0.68 (0.18) 0.90 (0.01) 0.08 (0.13) 0.47 (0.37)
f11 Discus 0.83 (0.01) 0.92 (0.01) 0.67 (0.20) 0.92 (0.00) 0.17 (0.16) 0.44 (0.34)
f12 BentCigar 0.55 (0.19) 0.87 (0.03) 0.49 (0.16) 0.73 (0.25) 0.08 (0.11) 0.41 (0.34)
f13 SharpRidge 0.71 (0.01) 0.91 (0.01) 0.62 (0.18) 0.88 (0.01) 0.17 (0.14) 0.44 (0.33)
f14 DifferentPowers 0.90 (0.01) 0.95 (0.01) 0.86 (0.04) 0.95 (0.00) 0.49 (0.18) 0.67 (0.28)
f15 RastriginRotated 0.25 (0.19) 0.49 (0.31) 0.14 (0.02) 0.49 (0.31) 0.10 (0.04) 0.13 (0.10)
f16 Weierstrass 0.47 (0.17) 0.87 (0.09) 0.27 (0.11) 0.72 (0.21) 0.17 (0.06) 0.26 (0.21)
f17 Schaffers10 0.71 (0.06) 0.90 (0.00) 0.55 (0.12) 0.90 (0.03) 0.29 (0.11) 0.43 (0.24)
f18 Schaffers1000 0.66 (0.08) 0.85 (0.00) 0.44 (0.11) 0.84 (0.09) 0.21 (0.08) 0.31 (0.18)
f19 GriewankRosenbrock 0.26 (0.15) 0.41 (0.22) 0.21 (0.01) 0.30 (0.02) 0.19 (0.03) 0.24 (0.06)
f20 Schwefel 0.68 (0.23) 0.46 (0.32) 0.44 (0.24) 0.19 (0.01) 0.25 (0.14) 0.18 (0.05)
f21 Gallagher101 0.73 (0.23) 0.95 (0.05) 0.47 (0.36) 0.44 (0.32) 0.29 (0.21) 0.36 (0.32)
f22 Gallagher21 0.72 (0.23) 0.81 (0.17) 0.50 (0.36) 0.62 (0.38) 0.26 (0.18) 0.32 (0.30)
f23 Katsuura 0.24 (0.08) 0.67 (0.22) 0.19 (0.01) 0.48 (0.11) 0.19 (0.01) 0.21 (0.11)
f24 LunacekBiRastrigin 0.12 (0.02) 0.14 (0.04) 0.10 (0.01) 0.12 (0.02) 0.08 (0.02) 0.09 (0.03)
@@
f8 Rosenbrock 0.38 (0.01) 0.41 (0.01) 0.37 (0.02) 0.40 (0.01) 0.24 (0.06) 0.33 (0.08)
f9 RosenbrockRotated 0.38 (0.01) 0.41 (0.00) 0.38 (0.01) 0.41 (0.00) 0.24 (0.06) 0.35 (0.07)
f10 EllipsoidRotated 0.22 (0.01) 0.29 (0.01) 0.19 (0.01) 0.28 (0.01) 0.14 (0.02) 0.20 (0.06)
@@
f12 BentCigar 0.34 (0.03) 0.51 (0.09) 0.34 (0.03) 0.45 (0.06) 0.05 (0.06) 0.24 (0.17)
f13 SharpRidge 0.42 (0.02) 0.51 (0.05) 0.42 (0.02) 0.49 (0.05) 0.31 (0.02) 0.39 (0.07)
f14 DifferentPowers 0.60 (0.01) 0.70 (0.01) 0.59 (0.01) 0.69 (0.00) 0.43 (0.04) 0.56 (0.10)
@@ 4.6. Structural Bias Analysis
Part of the IOH-xplainer framework is also our work on Structural Bias detection (biastoolbox, ; deepbias, ). The single-best and avg-best configurations are tested for structural bias and the results are given in Table 5 for Modular CMA-ES and Table 8 for Modular DE. To allow a visual in-depth analysis of the strength and type of structural bias, the toolbox provides for each test also a visualization of the distributions of locations of the final optima of a special test function $f_{0}$ (Kononova2015, ) (which should be uniform when no structural bias is present). It can be observed from the two tables that many single-best solutions exhibit structurally biased behaviour. For Modular DE, this is often bias near the boundaries of the search domain, while for Modular CMA-ES, it is often biased around the centre of the domain. It is important to take this structural bias effect into account, especially as we know that the optima of the BBOB benchmark are not uniformly distributed in the whole search domain (but in $[-4,4]^{d}$, and in the case of $f_{5}$ at the boundaries) (long2023bbob, ).
@@ 5. Additional Insights
In the proposed IOH-Xplainer pipeline, a lot of data is being collected about the behaviour and performance of algorithm configurations. Apart from the main insights we can derive using XAI methods on this data, there are additional benefits of storing and re-using this data. In this section we want to show a few ideas which can easily be done using the data collected and may prove beneficial to selecting, configuring and understanding these algorithm configurations.
The first “by-product” of our IOH-Xplainer toolbox, is that we can easily train machine learning models to predict a complete algorithm configuration (including module and hyper-parameter settings) based on landscape features. Basically performing automated algorithm configuration (and selection), similar to existing works such as (kerschke2019automated, ) and others (long2022learning, ; improvingnevergrad, ). Note however, that in this work we combine the selection and configuration of the algorithm as one machine learning task, using multi-output mixed classification and regression models. In our case, we limit the ML models to shallow decision trees and Random Forests for the sake of simplicity. To showcase this, we collected classical exploratory landscape analysis (ELA) features using a design of experiments of $1024$ samples on each BBOB instance used in our experiments. For each BBOB instance, we then took the single-best performing algorithm configuration (per modular framework), as the output of our machine learning pipeline. Using multi-output models, such as Random Forest and simple (and interpretable) decision trees, we can then train these models to predict good performing algorithm configurations based on the ELA features.
@@
Figure 8. Decision tree with max depth of 7 for elitist in modular CMA-ES with d=30.
As a second additional use of the collected data, we can build interpretable models such as shallow decision trees, to predict when to use a specific algorithm module. This types of “wizards” can provide a lot of additional insights on what kind of landscapes and what kind of landscape features are important to consider when selecting certain modules and settings. An example for the module elitist for ModCMA can be seen in Figure 8. In this example the mean mis-classification error (mmce) from the ELA level feature set seems to be most important for either seting elitist to true or false. Another example for the module crossover for the ModDE framework can be seen in Figure 9. All additional trees are available in our public Github repository  [^3] https://github.com/Basvanstein/ioh-explainer.
@@
Figure 9. Decision tree with max depth of 7 for crossover in modular DE with d=30.
@@ 6. Conclusion
In this work a novel framework for explainable benchmarking for iterative optimization heuristics is proposed and implemented as software tool called IOH-Xplainer. With explainable benchmarking, we mean to shift the benchmarking paradigm from very much performance focused (usually with a limited setup, biased towards the proposed algorithm (component)), towards a benchmarking practise to gain insights into the behaviour and influence of different modules and hyper-parameter settings. These insights can either be generated from existing algorithms, modular frameworks or newly developed operators or algorithms. The focus in explainable benchmarking lies in gaining understanding what search tactics and hyper-parameters contribute on what kind of black box problems and function landscapes. Using the IOH-Xplainer toolbox, practisioners and researchers in the field can now easily perform such a detailed and broad analysis of components with relative ease. The ability of IOH-Xplainer to provide detailed, explainable insights into algorithm performance marks a notable step forward in understanding and optimizing these complex systems. We showcase our proposed framework by two extensive use-cases, namely on the modular CMA-ES and the modular DE frameworks. Performing a detailed, explainable comparison of a staggering $52.128$ algorithm configurations over $24$ different benchmark functions ($5$ instances, $5$ random seeds) in both $5$ and $30$ dimensions. We analysed the collected data from these experiments with the help of explainable AI techniques and showed how this analysis can lead to new insights for both algorithm frameworks. Note, that the conclusions drawn for these frameworks are of course limited to the experimental setup, such as the hyper-parameter choices included and the upper limit of the chosen $10.000$ evaluation budget. Increasing the budget (or using different commonly used budgets) and adding additional hyper-parameter choices would be a set for future work to investigate.
In addition to the explanations and extensive analysis of these frameworks we also show that the data collected by the IOH-Xplainer toolbox can be used for additional tasks such as Automated Algorithm Configuration and explaining algorithm behaviour in relation with exploratory landscape analysis.
@@
We provide an open-source documented implementation of our package at (niki_van_stein_2024_10568760, ), including install and how-to-use guide. AOCC data collected from the runs are available as well, including detailed notebooks and scripts to repeat the experiments. All experiments are performed carbon neutral ($CO_{2}$-free) by using solar power.

### 23 Feb 2024 v2

@@ 4.5. Performance comparison between ModCMA and ModDE
While we have to stress that it is not the main aim of this work to compare ModCMA with ModDE, we can do a (limited) comparison in performance between the two modular frameworks using the data collected. In Table 9 we listed the performance of the single-best, avg-best and “default” performance of the frameworks in $5$ dimensions and in Table 10 for $30$ dimensions. The “default” performance denotes the average performance over all algorithm configurations. Results in boldface indicates a significant improvement (with p-value threshold of $0.05$) between the two frameworks in the same category (so single-best vs single-best, avg-best vs avg-best etc.). It can be observed that in general ModCMA seems to be better performing on most of the BBOB functions except for $f_{2},f_{3}$ and $f_{4}$ (all from the separable functions group). However, take into account that not all ModDE and not all ModCMA parameter options are tested (as this is computationally infeasible). If we look at the stability or tune-ability of the different frameworks we can conclude that in general ModDE gains more performance on average by optimizing the hyperparameters. For ModCMA in 30 dimensions, the gain (in terms of AOCC) to optimize for a specific function is on average not significant over taking the average best configuration (See Table 4).
@@
f1 Sphere 0.97 (0.00) 0.98 (0.00) 0.93 (0.00) 0.97 (0.00) 0.66 (0.26) 0.69 (0.30)
f2 Ellipsoid 0.94 (0.01) 0.91 (0.01) 0.88 (0.01) 0.90 (0.01) 0.48 (0.33) 0.39 (0.36)
f3 Rastrigin 0.89 (0.01) 0.38 (0.29) 0.66 (0.21) 0.30 (0.23) 0.29 (0.25) 0.12 (0.08)
f4 BuecheRastrigin 0.82 (0.03) 0.17 (0.16) 0.46 (0.27) 0.13 (0.02) 0.22 (0.20) 0.09 (0.04)
f5 LinearSlope 1.00 (0.00) 1.00 (0.00) 0.99 (0.01) 0.99 (0.00) 0.94 (0.17) 0.95 (0.16)
f6 AttractiveSector 0.88 (0.05) 0.95 (0.01) 0.75 (0.26) 0.90 (0.01) 0.30 (0.22) 0.48 (0.34)
f7 StepEllipsoid 0.91 (0.01) 0.95 (0.00) 0.89 (0.02) 0.94 (0.03) 0.37 (0.25) 0.50 (0.31)
f8 Rosenbrock 0.81 (0.06) 0.91 (0.02) 0.71 (0.23) 0.85 (0.02) 0.25 (0.21) 0.42 (0.34)
f9 RosenbrockRotated 0.80 (0.03) 0.92 (0.01) 0.70 (0.24) 0.86 (0.01) 0.23 (0.21) 0.45 (0.33)
f10 EllipsoidRotated 0.78 (0.03) 0.90 (0.01) 0.68 (0.18) 0.89 (0.01) 0.08 (0.13) 0.39 (0.36)
f11 Discus 0.83 (0.01) 0.92 (0.01) 0.67 (0.20) 0.91 (0.01) 0.17 (0.16) 0.37 (0.32)
f12 BentCigar 0.55 (0.19) 0.86 (0.07) 0.49 (0.16) 0.73 (0.23) 0.08 (0.11) 0.33 (0.33)
f13 SharpRidge 0.71 (0.01) 0.90 (0.01) 0.62 (0.18) 0.87 (0.01) 0.17 (0.14) 0.37 (0.31)
f14 DifferentPowers 0.90 (0.01) 0.95 (0.01) 0.86 (0.04) 0.95 (0.00) 0.49 (0.18) 0.62 (0.27)
f15 RastriginRotated 0.25 (0.19) 0.48 (0.02) 0.14 (0.02) 0.43 (0.19) 0.10 (0.04) 0.13 (0.10)
f16 Weierstrass 0.47 (0.17) 0.81 (0.16) 0.27 (0.11) 0.69 (0.19) 0.17 (0.06) 0.26 (0.19)
f17 Schaffers10 0.71 (0.06) 0.90 (0.02) 0.55 (0.12) 0.87 (0.05) 0.29 (0.11) 0.41 (0.21)
f18 Schaffers1000 0.66 (0.08) 0.82 (0.12) 0.44 (0.11) 0.68 (0.16) 0.21 (0.08) 0.30 (0.16)
f19 GriewankRosenbrock 0.26 (0.15) 0.40 (0.23) 0.21 (0.01) 0.31 (0.02) 0.19 (0.03) 0.24 (0.06)
f20 Schwefel 0.68 (0.23) 0.36 (0.11) 0.44 (0.24) 0.21 (0.07) 0.25 (0.14) 0.19 (0.05)
f21 Gallagher101 0.73 (0.23) 0.84 (0.23) 0.47 (0.36) 0.53 (0.32) 0.29 (0.21) 0.40 (0.31)
f22 Gallagher21 0.72 (0.23) 0.79 (0.26) 0.50 (0.36) 0.46 (0.33) 0.26 (0.18) 0.36 (0.30)
f23 Katsuura 0.24 (0.08) 0.66 (0.22) 0.19 (0.01) 0.49 (0.20) 0.19 (0.01) 0.21 (0.10)
f24 LunacekBiRastrigin 0.12 (0.02) 0.14 (0.03) 0.10 (0.01) 0.13 (0.02) 0.08 (0.02) 0.09 (0.02)
@@
f8 Rosenbrock 0.38 (0.01) 0.41 (0.01) 0.37 (0.02) 0.40 (0.01) 0.24 (0.06) 0.32 (0.08)
f9 RosenbrockRotated 0.38 (0.01) 0.41 (0.00) 0.38 (0.01) 0.41 (0.00) 0.24 (0.06) 0.34 (0.06)
f10 EllipsoidRotated 0.22 (0.01) 0.29 (0.01) 0.19 (0.01) 0.28 (0.01) 0.14 (0.02) 0.19 (0.06)
@@
f12 BentCigar 0.34 (0.03) 0.51 (0.09) 0.34 (0.03) 0.45 (0.06) 0.05 (0.06) 0.21 (0.17)
f13 SharpRidge 0.42 (0.02) 0.51 (0.05) 0.42 (0.02) 0.49 (0.05) 0.31 (0.02) 0.38 (0.07)
f14 DifferentPowers 0.60 (0.01) 0.70 (0.01) 0.59 (0.01) 0.69 (0.00) 0.43 (0.04) 0.54 (0.10)
@@ 4.6. Structural Bias Analysis
Part of the IOHxplainer framework is also the work on Structural Bias detection (biastoolbox, ; deepbias, ). The single-best and avg-best configurations are tested for structural bias and the results of the combined statistical and deep-learning SB detection methods are given in Table 5 for Modular CMA-ES and Table 8 for Modular DE.
To allow a visual in-depth analysis of the strength and type of structural bias, the toolbox provides for each test also a visualization of the distributions of locations of the final optima of a special test function $f_{0}$ (Kononova2015, ) (which should be uniform when no structural bias is present). Each of the reported SB cases has been verified and if needed corrected by visual inspection of these distributions. This correction is particularly applied when the detection method is predicting the gaps/cluster class, since we know that centre bias is often misclassified as gaps/clusters (in theory centre bias is a special version of cluster bias). All visualizations are available in our Github repository (code_xplainer_niki_van_stein_2024_10568760, ).
It can be observed from the two tables that many single-best solutions exhibit structurally biased behaviour. For Modular DE, this is often bias near the boundaries of the search domain, while for Modular CMA-ES, it is almost always biased around the centre of the domain except for a few configurations 30 dimensions. It is important to take this structural bias effect into account, especially as we know that the optima of the BBOB benchmark are not uniformly distributed in the whole search domain (but in $[-4,4]^{d}$, and in the case of $f_{5}$ – at the boundaries) (long2023bbob, ), which might explain why CMA-ES is often outperforming on this benchmark.
@@ 5. Additional Insights
In the proposed IOHxplainer pipeline, a lot of data is being collected about the behaviour and performance of algorithm configurations. Apart from the main insights we can derive using XAI methods on this data, there are additional benefits of storing and re-using this data. In this section we want to show a few ideas which can easily be done using the data collected and may prove beneficial to selecting, configuring and understanding these algorithm configurations.
The first “by-product” of our IOHxplainer toolbox, is that we can easily train machine learning models to predict a complete algorithm configuration (including module and hyper-parameter settings) based on landscape features. Basically performing automated algorithm configuration (and selection), similar to existing works such as (kerschke2019automated, ) and others (long2022learning, ; improvingnevergrad, ). Note however, that in this work we combine the selection and configuration of the algorithm as one machine learning task, using multi-output mixed classification and regression models. In our case, we limit the ML models to shallow decision trees and Random Forests for the sake of simplicity. To showcase this, we collected classical exploratory landscape analysis (ELA) features using a design of experiments of $1024$ samples on each BBOB instance used in our experiments. For each BBOB instance, we then took the single-best performing algorithm configuration (per modular framework), as the output of our machine learning pipeline. Using multi-output models, such as Random Forest and simple (and interpretable) decision trees, we can then train these models to predict good performing algorithm configurations based on the ELA features.
@@
Figure 8. Decision tree with max depth of 7 for elitist in modular CMA-ES with d=30. Each node of the tree represent a split where the distribution of the feature used for the split is visualized. Colors in the distribution denote the different labels (in this case yellow is ‘false’ and green is ‘true’ for the elitist module). The black arrow below the distribution denotes the splitting threshold.
As a second additional use of the collected data, we can build interpretable models such as shallow decision trees, to predict when to use a specific algorithm module. This types of “wizards” can provide a lot of additional insights on what kind of landscapes and what kind of landscape features are important to consider when selecting certain modules and settings. An example for the module elitist for ModCMA can be seen in Figure 8. In this example the mean mis-classification error (mmce) from the ELA level feature set seems to be most important for either seting elitist to true or false. Another example for the module crossover for the ModDE framework can be seen in Figure 9. All additional trees are available in our public Github repository (code_xplainer_niki_van_stein_2024_10568760, ).
@@
Figure 9. Decision tree with max depth of 7 for crossover in modular DE with d=30. Each node of the tree represent a split where the distribution of the feature used for the split is visualized. Colors in the distribution denote the different labels (in this case yellow is ‘bin’ and green is ‘exp’ crossover). The black arrow below the distribution denotes the splitting threshold.
@@ 6. Conclusion
In this work a novel framework for explainable benchmarking for iterative optimization heuristics is proposed and implemented as software tool called IOHxplainer. With explainable benchmarking, we mean to shift the benchmarking paradigm from very much performance focused (usually with a limited setup, biased towards the proposed algorithm (component)), towards a benchmarking practise to gain insights into the behaviour and influence of different modules and hyper-parameter settings. These insights can either be generated from existing algorithms, modular frameworks or newly developed operators or algorithms. The focus in explainable benchmarking lies in gaining understanding what search tactics and hyperparameters contribute on what kind of black box problems and function landscapes. Using the IOHxplainer toolbox, practisioners and researchers in the field can now easily perform such a detailed and broad analysis of components with relative ease. The ability of IOHxplainer to provide detailed, explainable insights into algorithm performance marks a notable step forward in understanding and optimizing these complex systems. We showcase our proposed framework by two extensive use-cases, namely on the modular CMA-ES and the modular DE frameworks. Performing a detailed, explainable comparison of a staggering $52.128$ algorithm configurations over $24$ different benchmark functions ($5$ instances, $5$ random seeds) in both $5$ and $30$ dimensions. We analysed the collected data from these experiments with the help of explainable AI techniques and showed how this analysis can lead to new insights for both algorithm frameworks. Note, that the conclusions drawn for these frameworks are of course limited to the experimental setup, such as the hyper-parameter choices included and the upper limit of the chosen $10.000$ evaluation budget. Increasing the budget (or using different commonly used budgets) and adding additional hyper-parameter choices would be a set for future work to investigate.
In addition to the explanations and extensive analysis of these frameworks we also show that the data collected by the IOHxplainer toolbox can be used for additional tasks such as Automated Algorithm Configuration and explaining algorithm behaviour in relation with exploratory landscape analysis.
@@
We provide an open-source documented implementation of our package at (code_xplainer_niki_van_stein_2024_10568760, ), including install and how-to-use guide. AOCC data collected from the runs are available as well, including detailed notebooks and scripts to repeat the experiments. All experiments are performed carbon neutral ($CO_{2}$-free) by using solar power.